Can you please help me solve this trigonometric identities question: 1)Prove the following identities.
a) 1-cos^2 x=tan^2 xcos^2 x

Answer by vleith(2983)

1-cos^2 x=tan^2xcos^2 x
1-cos^2 x=(tan^2x)cos^2 x
1-cos^2 x=(sin^2x/cos^2x)cos^2 x
1-cos^2 x=sin^2x
1 = sin^2 + cos^2
You know the last one is true. So just work backwards from there